### v2

# LLM4GEN: Leveraging Semantic Representation of LLMs for Text-to-Image Generation

Abstract
Diffusion models have exhibited substantial success in text-to-image generation. However, they often encounter challenges when dealing with complex and dense prompts involving multiple objects, attribute binding, and long descriptions. In this paper, we propose a novel framework called LLM4GEN, which enhances the semantic understanding of text-to-image diffusion models by leveraging the representation of Large Language Models (LLMs). It can be seamlessly incorporated into various diffusion models as a plug-and-play component. A specially designed Cross-Adapter Module (CAM) integrates the original text features of text-to-image models with LLM features, thereby enhancing text-to-image generation. Additionally, to facilitate and correct entity-attribute relationships in text prompts, we develop an entity-guided regularization loss to further improve generation performance. We also introduce DensePrompts, which contains $7,000$ dense prompts to provide a comprehensive evaluation for the text-to-image generation task. Experiments indicate that LLM4GEN significantly improves the semantic alignment of SD1.5 and SDXL, demonstrating increases of 9.69% and 12.90% in color on T2I-CompBench, respectively. Moreover, it surpasses existing models in terms of sample quality, image-text alignment, and human evaluation.
Introduction
Recently, diffusion models (Song et al. 2020; Rombach et al. 2022; Chen et al. 2023a) have made significant progress in text-to-image (T2I) generation models, such as Imagen (Saharia et al. 2022), DALL-E (Betker et al. 2023), and Stable Diffusion (Rombach et al. 2022; Podell et al. 2023). However, they often encounter challenges in generating images given complex and dense prompt descriptions, such as attribute binding and multiple objects (Huang et al. 2023).
Figure 1: Architecture comparison between (a) LLM-guidance models (b) LLM-alignment models and (c) our proposed LLM4GEN.
With the emergence of powerful linguistic representations from Large Language Models (LLMs), there has been an increasing trend in leveraging LLMs to aid in T2I generation. Current methods mainly consist of two categories: LLM-guidance models (Yang et al. 2024; Feng et al. 2022) and LLM-alignment models (Wu et al. 2023; Hu et al. 2024; sd3; Zhao et al. 2024). LLM-guidance models harness the reasoning capability of LLMs and the Layout model to generate controllable images, as illustrated in Fig. 1 (a). However, these methods require separating LLMs from external models, resulting in redundancy in both inference time and the overall framework. While LLM-alignment models utilize LLMs to exploit the representational capacity, they demand substantial training data to align LLM representations with the diffusion model, as shown in Fig. 1 (b).
To address aforementioned challenges, we propose LLM4GEN, a novel framework that implicitly leverages the powerful semantic representations of LLMs to enhance the original text encoder for T2I GENeration, as illustrated in Fig. 1 (c). Specifically, we design an efficient Cross-Adapter Module (CAM) to implicitly integrate the semantic representation of LLMs with original text encoders that have limited representational capabilities, such as CLIP text encoder (Radford et al. 2021). We apply cross-attention to the representations of both encoder-only LLMs (e.g., Llama (Zhang et al. 2024)) and decoder-only LLMs (e.g., T5 (Raffel et al. 2020)) alongside CLIP text embeddings, and then concatenate the fused embedding with the original CLIP text embedding. This CAM module significantly enhances the performance of T2I diffusion models while preserving the original text encoder representations, thereby reducing the need for extensive training data. Additionally, we introduce an entity-guidance regularization loss that penalizes mismatches between the activation maps of entities and their corresponding attributes in the text, improving the model’s ability to accurately comprehend and represent the main subjects in the generated images. As evidenced in Fig. 2, our proposed method exhibits strong performance in T2I generation.
Figure 2: Image generation using concise and dense prompts, with colored text highlighting key entities or attributes(Zoom in for details).
To comprehensively assess the image generation capabilities of T2I models, we develop a comprehensive benchmark named DensePrompts, an extension of T2I-CompBench (Huang et al. 2023), which incorporates over 7,000 compositional prompts. The construction of this benchmark involves leveraging LLMs for complex text descriptions, followed by manual refinement. Results from performance metrics and human evaluations consistently demonstrate that LLM4GEN’s representational capability surpasses other existing methods.
Overall, our contributions are as follows:
  • We propose a novel framework that leverages the powerful representational capabilities of LLMs to assist in text-to-image (T2I) generation. Specifically, we design a Cross-Adapter to integrate LLM representations and introduce an entity-guidance regularization loss to enhance semantic understanding.
  • To assess performance with long-text prompts, we introduce DensePrompts, a benchmark designed to evaluate both aesthetic quality and image-text alignment.
  • Our designed LLM4GEN can be seamlessly integrated into existing diffusion models like SD1.5 (Rombach et al. 2022) and SDXL (Podell et al. 2023). Experiments show that LLM4GEN exhibits superior performance in sample quality, image-text alignment, and human evaluation compared with existing state-of-the-art models.
Related Work
Large Language Models
Large language models (LLMs) (Chang et al. 2023) have shown powerful generalization ability in various NLP tasks. Recent LLMs, e.g., GPTs (Brown et al. 2020), LLaMA (Touvron et al. 2023), OPT (Zhang et al. 2022), PaLM (Chowdhery et al. 2022) are all equipped with billions of parameters, enabling the intriguing capability for in-context learning and demonstrating excellent zero-shot performance across various tasks. Certain Multi-modal LLMs (MLLMs) (Achiam et al. 2023; Zhu et al. 2023; Bai et al. 2023) have integrated visual and audio modalities, enhancing intelligent interactions with the help of LLMs. (Pang et al. 2024) shows that the frozen LLMs can further integrate visual understanding. Recent works (Lian et al. 2024; Sun et al. 2024; Yang et al. 2024) use LLMs to create improved text prompts or bounding box layouts for high-quality text-to-image generation. However, these existing works only consider LLMs as simple condition generators, e.g., text prompts or layout planning. In this paper, we harness the representation capabilities of LLMs to enhance text-to-image generation, emphasizing their significant representational power beyond simple text output.
Text-to-Image Diffusion Models Text-to-image generation aims to create images with given prompts. Diffusion models (Song and Ermon 2019, 2020; Song et al. 2020) have demonstrated remarkable performance in image generation. These models use added Gaussian noise for a forward process and can generate diverse, high-quality images through an inverse process from random Gaussian noise. GLIDE (Nichol et al. 2022) utilizes CLIP (Radford et al. 2021) text encoder to enhance the image-text alignment. Latent Diffusion Models (LDMs) (Rombach et al. 2022) transfer the diffusion process from pixel to latent space. Recent models such as SD-XL (Podell et al. 2023), DALL-E 3 (Betker et al. 2023), and Dreambooth (Ruiz et al. 2023) have significantly enhanced image quality and text-image alignment using various perspectives, such as training strategies and scaling training data. Despite these notable advancements, generating high-fidelity images aligned with complex textual prompts remains challenging. In this paper, we propose LLM4GEN, which leverages the robust representation capabilities of LLMs to facilitate image generation from textual descriptions.
Figure 3: The overview of LLM4GEN. (a) Framework. (b) Cross-Adapter Module.
Methodology
LLM4GEN
Framework
The proposed LLM4GEN, which contains a Cross-Adapter Module (CAM) and the UNet, is illustrated in Fig. 3 (a). In this paper, we explore stable diffusion (Rombach et al. 2022; Podell et al. 2023) as the base text-to-image diffusion model, and the vanilla text encoder is from CLIP (Radford et al. 2021). LLM4GEN leverages the strong capability of LLMs to assist in text-to-image generation. The CAM extracts the representation of a given prompt via the combination of LLM and CLIP text encoder. The fused text embedding is enhanced by leveraging the pre-trained knowledge of LLMs through the simple yet effective CAM. By feeding the fused text embedding, LLM4GEN iteratively denoises the latent vectors with the UNet and decodes the final vector into an image with the VAE.
Cross-Adapter Module
The CAM connects the LLMs and the CLIP text encoder using a cross-attention layer, followed by concatenation with the representation of the CLIP text encoder. The last hidden state of the LLMs is extracted as LLMs feature $c_{l}$. The feature of CLIP text encoder is denoted as $c_{t}$, and we perform a cross-attention to fuse them:
$Q=W_{q}(c_{l}),K=W_{k}(c_{t}),V=W_{v}(c_{t})$ (1)
$c_{l}^{\prime}=\operatorname{CrossAttention}(Q,K,V)=\operatorname{softmax}\left(Q\cdot K^{T}\right)\cdot V$ (2)
where $W_{q}$, $W_{k}$, $W_{v}$ are the trainable linear projection layers. The output embedding dimension is the same as that of CLIP text encoder. Then the final fused text embedding of the CAM is:
$x=\operatorname{CA}(x,\operatorname{Concat}(\lambda\cdot c_{l}^{\prime},c_{t}))=\lambda\cdot\operatorname{CA}(x,c_{l})+\operatorname{CA}(x,c_{t})$ (3)
where $\operatorname{Concat}$ denotes concatenation in the sequence dimension, and $\lambda$ is the balance factor, $x$ denotes the latent noise, $\operatorname{CA}$ is the cross-attention module within the UNet module. Overall, our designed Cross-Adapter Module implicitly facilitates the strong representation of LLMs with a residual fusion manner, without utilizing extensive training data and resources to condition the latent vectors on text embeddings. Notably, our LLM4GEN is compatible with both decoder-only and encoder-only LLMs and we evaluate on Llama-2 7B/13B (Touvron et al. 2023) and T5-XL (Brown et al. 2020) in further experiments.
Figure 4: Statistic of DensePrompts benchmark compared with other benchmarks.
Entity-Guidance Regularization Loss
Current text-to-image generation models often encounter confusion and omissions when generating multiple entities. We utilize a parser to analyze the prompt $\mathcal{P}$, extracting a set of attribute-entity pairs $\mathcal{S}=\sum_{i=1}^{N}\{a_{i},e_{i}\}$, where $e_{i}$ and $a_{i}$ represent the entity name and its corresponding attribute, respectively, and $N$ denotes the number of parsed pairs. Subsequently, we can calculate the active similarity map as:
$\mathcal{A}_{i}=\operatorname{softmax}\left(\frac{QK_{i}^{T}}{\sqrt{d}}\right)$ (4)
where the query $Q$ is derived from the latent representation, the key $K_{i}$ is derived from the token embedding of $p$, and $d$ is the latent dimension. $\mathcal{A}_{a}$ and $\mathcal{A}_{o}$ indicate the similarity maps for the attribute word and the entity, respectively. Subsequently, we impose a penalty on these similarity maps on all UNet layers as:
$\mathcal{L}_{reg}=\frac{1}{N\cdot L}\sum_{i=1}^{N}\sum_{l=1}^{L}||\mathcal{A}_{a}^{i-l}-\mathcal{A}_{o}^{i-l}||^{2}$ (5)
where |||||$\cdot$||| | ⋅ | | represents the L2 distance, $L$ is the layer numbers.
Overall, based on the framework described above, the training loss of LLM4GEN is formulated as:
$\mathcal{L}=\mathbb{E}_{\varepsilon(x),\epsilon\sim\mathcal{N}(0,1),t}[\parallel\epsilon-\epsilon_{\theta}(z_{t},t)\parallel]_{2}^{2}]+\alpha\cdot\mathcal{L}_{reg}$ (6)
where $z_{t}$ can be obtained from the encoder $\mathcal{E}$, ans latent vectors from $p(z)$ can be decoded to images through the decoder $\mathcal{D}$. In this paper, we address the limited representation of CLIP as a text encoder by leveraging the capabilities of large language models (LLMs) to enhance the text encoder of the LDMs.
DensePrompts Benchmark
A comprehensive benchmark is crucial for evaluating the image-text alignment of generated images. Current benchmarks, e.g., MSCOCO (Lin et al. 2014) and T2I-CompBench (Huang et al. 2023), primarily consist of concise textual descriptions, are not comprehensive enough to describe a diverse range of objects. Thus, we introduce a new comprehensive and complicated benchmark called DensePrompts, comprising lengthy textual descriptions.
Initially, we collect 100 images from the Internet, comprising 50 real and 50 generated images, each with intricate details. Leveraging the robust image comprehension capabilities of GPT-4V (OpenAI 2023), we utilize it to provide detailed descriptions for these 100 images, encompassing object attributes and their relationships, thereby generating comprehensive prompts abundant in semantic details. We employ GPT-4 (Achiam et al. 2023) to produce massive long texts based on generated prompts mentioned above. DensePrompts provides more than 7,000 extensive prompts whose average word length is more than 40. Word statistics of DensePrompts are outlined in Fig. 4. To assess the performance, DensePrompts benchmark incorporates CLIP Score (Radford et al. 2021) and Aesthetic Score (Schuhmann ). Combining our proposed DensePrompts with T2I-CompBench, we establish a comprehensive evaluation in text-to-image generation.
Experiments
Experimental Details
Framework and Implementation Details
In this paper, we explore LLM4GEN based on SD1.5 and SDXL, denoted as LLM4GENSD1.5 and LLM4GENSDXL. We utilize T5-XL and CLIP text encoder (CLIP ViT-L/14) as the text tower. The sequence length of the LLMs is set to $128$. We use 10M text-image pairs collected from LAION-2B(Schuhmann et al. 2021) and Internet. Training is conducted on 8NVIDIA A100 GPUs with the learning rates of 2e-5 and 1e-5 for LLM4GENSD1.5 and LLM4GENSDXL, respectively. The batch size is set to 256 and 128. The training steps are set to 20k and 40k. Additionally, we further train LLM4GENSDXL using 2M high-quality data with 1024 resolution. During inference, we utilize DDIM sampler (Song, Meng, and Ermon 2020) for sampling with 50 steps and the classifier free guidance scale to 7.5.
Evaluation Benchmarks
We comprehensively evaluate proposed LLM4GEN via four primary benchmarks, e.g. MSCOCO (Lin et al. 2014), T2I-CompBench (Huang et al. 2023), our proposed DensePrompts benchmark, and User Study.
[…]
Color $\uparrow$ Shape $\uparrow$ Texture $\uparrow$ Spatial $\uparrow$ Non-Spatial $\uparrow$
Composable Diffusion (Liu et al. 2022) 40.63 32.99 36.45 8.00 29.80 28.98
Structured Diffusion (Feng et al. 2022) 49.90 42.18 49.00 13.86 31.11 33.55
Attn-Exct v2 (Chefer et al. 2023) 64.00 45.17 59.63 14.55 31.09 34.01
GORS (Huang et al. 2023) 66.03 47.85 62.87 18.15 31.93 33.28
DALL-E 2 (Ramesh et al. 2022) 57.50 54.64 63.74 12.83 30.43 36.96
PixArt-$\alpha$ (Chen et al. 2023b) 68.86 55.82 70.44 20.82 31.79 41.17
SUR-Adapter* (Zhong et al. 2023) 38.93 36.73 39.21 13.21 30.81 31.43
ELLASDXL (Hu et al. 2024) 72.60 56.34 66.86 22.14 30.69 -
SD1.5 (Rombach et al. 2022) 37.65 35.76 41.56 12.46 30.79 30.80
LLM4GENSD1.5 (Ours) 47.34 45.35 55.39 14.79 31.00 33.97
$\Delta$ (Margin) +9.69 +9.59 +13.83 +2.33 +0.21 +3.17
SDXL (Podell et al. 2023) 63.69 54.08 56.37 20.32 31.10 40.91
LLM4GENSDXL (Ours) 76.59 59.24 70.86 24.12 32.10 42.19
$\Delta$ (Margin) +12.90 +5.16 +14.49 +3.80 +1.00 +1.28
Table 1: Evaluation results (%) on T2I-CompBench (Huang et al. 2023). The higher is better, and the best results are highlighted in bold. * denotes that we calculate the metrics using the official weights and code provided.
[…]
SD1.5 (Rombach et al. 2022) 26.89 32.24 28.66
SD1.5 (ft) 25.48 33.53 29.10
LLM4GENSD1.5 25.20 34.24 29.45
SDXL (Podell et al. 2023) 24.75 34.91 30.10
LLM4GENSDXL 24.21 35.10 30.91
Table 2: Quantitative comparison on text-to-image generation models on the subset of MSCOCO (Lin et al. 2014) dataset.
Figure 5: Aesthetic Score and CLIP Score (%) on DensePrompts benchmark.
Figure 6: Results on user study regarding the sample quality and image-text alignment of different models.
Figure 7: A comparative analysis of LLM4GEN and other state-of-the-art diffusion models using PartiPrompts (Yu et al. 2022) and our proposed DensePrompts as prompts. The last row represents the prompts used.
Performance Comparisons and Analysis
Fidelity assessment on MSCOCO benchmark Experimental results on MSCOCO benchmark are shown in Tab. 2. LLM4GEN notably enhances the sample quality and image-text alignment, resulting in improvements of 1.79 and 0.54 on FID compared to SD1.5 and SDXL, respectively. Furthermore, we assess the performance of SD1.5 after extensive fine-tuning with the same training dataset. This modified version, SD1.5 (ft), surpasses the original SD1.5, yet LLM4GENSD1.5 still exhibits superior performance over SD1.5 (ft). This underscores the potent representation of our proposed LLM4GEN and its contribution to text-to-image generation.
Evaluation on T2I-CompBench For T2I-CompBench comparison, we select the recent text-to-image generative models for comparison, e.g., Composable Diffusion, Structured Diffusion, Attn-Exct v2, GORS, DALLE 2, PixArt-$\alpha$, ELLASDXL, SD1.5, and SDXL. Experimental results shown in Tab. 1 demonstrate the distinctive performance of LLM4GENSDXL in T2I-CompBench evaluation, underlining its advancements in attribute binding, object relationship, and mastery in rendering complex compositions. LLM4GEN shows considerable improvement in color, shape, and texture, showcasing enhancements up to +12.90% in color, +5.16% in shape, and +14.49% in texture with SDXL, respectively. LLM4GENSDXL also marks considerable progress in both spatial and non-spatial evaluations, with 3.80% and 1.00% lift, respectively. Furthermore, when compared with PixArt-$\alpha$, which employs T5-XL as its text encoder, LLM4GENSDXL surpasses it in several aspects, such as a notable 7.73% lead in color metric. Moreover, LLM4GENSDXL outperforms ELLASDXL. These results verify the potent synergy of LLMs representations in augmenting the sample quality and image-text alignment of diffusion models.
Evaluation on DensePrompts We compare our LLM4GEN with PixArt-$\alpha$, Playground v2, SD1.5, and SDXL on our DensePrompts benchmark. As shown in Fig. 5, LLM4GENSDXL achieves the highest Aesthetic Score and CLIP Score among these models. PixArt-$\alpha$ outperforms SDXL due to its T5-XL text encoder for dense prompts. LLM4GEN excels in understanding and interpreting dense prompts, resulting in high-quality images with strong image-text alignment. This performance is attributed to the powerful representation of LLMs and the effective adaptation of the original CLIP text encoder via our CrossAdapter Module. We attribute this performance to th powerful representation of LLMs and the effective adaptation of the original CLIP text encoder via our CrossAdapter Module.
Quantitive Results. To thoroughly evaluate our proposed LLM4GEN framework, we present the qualitative results on the short prompts provided by PartiPrompts (Yu et al. 2022) in the first 4 columns and on the dense prompts provided by DensePrompts in the last 3 columns in Fig. 7. The results indicate that our proposed LLM4GENSD1.5 and LLM4GENSDXL exhibit strong text-image alignment and superior dense prompt generation compared to the recent PixArt-$\alpha$, especially in handling the multiple objects and attribute binding.
User Study We conduct the user study on various combinations of existing methods and LLM4GENSDXL. For each pairing, we assess two criteria: sample quality and image-text alignment. Users are tasked with evaluating the aesthetic appeal and semantic understanding of images with identical text to determine the superior one based on these assessment criteria. Subsequently, we compute the percentage scores for each model, as shown in Fig. 6. The results showcase our LLM4GENSDXL exhibits comparative advantages over both SD1.5 and SDXL. Specifically, LLM4GENSDXL achieves 60.3% and 66.5% higher voting preferences compared to SDXL in terms of Aesthetic and Semantic, respectively. Notably, LLM4GENSDXL also delivers competitive results when compared to DALL-E 3.
Ablation Studies
Impact of Cross-Adapter Module
Due to limited computing sources, we evaluate the impact of various architectural enhancements on SD1.5, as outlined in Tab. 4. Our configurations explore different methods for integrating LLMs embeddings: (1) the baseline SD1.5 model, (2) SD1.5 finetune-CLIP, the result of fine-tuning the original text-encoder of SD1.5, (3) MLP or CrossAttention, which utilizes a simple linear layer or cross-attention layer to transform LLM embeddings, (4) MLP + Concat, representing a process where LLMs embeddings are projected to the same dimension as the original text embeddings before concatenation, (5) CrossAttention + Concat, (6) CLIP as Q and LLM as KV, refering to converting the position of Q and KV in CAM. Results show that configuration (2) indeed brings an improvement of the original SD1.5, yet, due to the limited semantic representation of CLIP, the results still remain subpar. Interestingly, simply concatenating the original text embeddings (configuration 3 & 4) provides a significant boost over base SD1.5. This suggests that direct representation alignment between LLMs and the latent vector is challenging, and enhancing the original text embeddings with LLM embeddings is sufficient to improve image-text alignment. In our LLM4GEN, the LLM representation is employed as Q, while the original text encoder serves as K and V. We also examine the impact of rearranging the position of Q and KV in the CAM module. The results, as demonstrated in configuration 6, indicate that our LLM4GEN (configuration 7) exceeds it, showcasing a 3.65% enhancement in color. This emphasizes the substantial benefits of incorporating our Cross-Adapter Module to enrich the representation of the original text encoder and the image-text alignment of generated images.
LLMs Attribute Binding
Color $\uparrow$ Shape $\uparrow$ Texture $\uparrow$
SD1.5 (Rombach et al. 2022) 37.65 35.76 41.56
Llama-2/7B (Touvron et al. 2023) 43.21 40.12 48.91
Llama-2/13B (Touvron et al. 2023) 44.98 41.03 49.21
T5-XL (Raffel et al. 2020) 47.34 45.35 55.39
Table 3: Impact (%) of Different LLMs based on SD1.5.
Module Attribute Binding
Color $\uparrow$ Shape $\uparrow$ Texture $\uparrow$
(1)  SD1.5 37.65 35.76 41.56
(2)  SD1.5 finetune-CLIP 38.76 36.85 42.56
(3)  MLP or CrossAttention 39.25 37.68 42.89
(4)  MLP + Concat 40.24 38.23 44.39
(5)  CrossAttention + Concat 42.31 39.43 46.21
(6)  CLIP as Q and LLM as KV 43.69 42.82 49.56
(7)  Ours 47.34 45.35 55.39
Table 4: Impact (%) of the designed Cross-Adapter Module.
Impact of Different LLMs
The analysis encompasses a comparative evaluation between base SD1.5 and the enhancements achieved through the integration of Llama-2/7B, Llama-2/13B, and T5-XL. As depicted in Tab. 3, the inclusion of any LLM improves upon the performance of SD1.5. Importantly, Llama-v2/13B outperforms Llama-v2/7B, demonstrating that LLMs with greater capacity excel in extracting more nuanced semantic embeddings. Furthermore, when compared to decoder-only LLMs, T5-XL encoder demonstrates advantages in semantic comprehension, confirming its superior suitability for enhancing text-to-image generation.
Further Analysis
Method #Images (M $\downarrow$) #GPU days ($\downarrow$) Performance ($\uparrow$)
PixArt-$\alpha$ (Chen et al. 2023b) 25 753 68.86
ParaDiffusion (Wu et al. 2023) 500 392 -
ELLASDXL (Hu et al. 2024) 30 112 72.60
LLM4GENSDXL 10 50 76.59
Table 5: Training resources comparison, including the scale of training data and computing cost.
Figure 8: Training data scaling analysis.
Figure 9: Cross-attention visualization (Tang et al. 2022) for two generated images. The two rows are SD1.5 and LLM4GENSD1.5, respectively.
Scaling Analysis
As illustrated in Fig. 8, we conduct an extensive analysis of the scalability of our proposed LLM4GEN model with respect to training data. The results conclusively demonstrate that as the quantity of training data increases, the performance of our model exhibits consistent and significant growth, thereby confirming its scalability. However, increasing the dataset scale from 5M to 10M resulted in minimal performance improvement on the generated images. Consequently, we use 10M text-image pairs for training our LLM4GEN.
Training Efficiency
When evaluating the effectiveness of integrating LLMs into text-to-image generation models, LLM4GENSDXL stands out for its remarkable efficiency and performance. LLM4GEN achieves significant reductions in both training data requirements and computational costs. It utilizes only 10 million data, a 66% reduction compared to ELLA, and demands merely 50 GPU days for training, drastically lower than PixArt-$\alpha$ (25 million data, 753 GPU days) and ParaDiffusion (500 million data, 392 GPU days). Despite this, LLM4GENSDXL achieves a superior color metric performance of 73.29%. This notable difference underscores LLM4GEN’s ability to substantially reduce both training data and computational costs while establishing a new standard for performance efficiency.
Cross-attention Visualization.
Fig. 9 shows the cross-attention visualization of SD1.5 and LLM4GENSD1.5, respectively. The heatmaps reveal that our proposed LLM4GEN method demonstrates a superior ability to capture relationships between attributes, such as ”blue” and ”sheep,” as illustrated in Fig. 9 (a). We attribute this enhanced capability to the increased semantic richness afforded by the robust representations of LLMs.
More visualization and experimental results are shown in the Appendix.
Conclusion
In this paper, we propose LLM4GEN, an end-to-end text-to-image generation framework. Specifically, we design an efficient Cross-Adapter Module to leverage the powerful representation of LLMs, thereby enhancing the original text representation of diffusion models. Despite using fewer training data and computational resources, LLM4GEN outperforms current state-of-the-art text-to-image diffusion models in sample quality and image-text alignment. To optimize consistency in entity-attribute relationships of generated images, we design an entity-guided regularization loss. Additionally, we introduce the DensePrompts benchmark to promote the generation of images with dense information and provide a comprehensive evaluation framework. Extensive experiments have shown that our proposed method achieves competitive performance.